### v2

@@ 1 Introduction
Theorem 1 characterizes whether each top-$k$ equilibrium exists. This family of equilibria has an intuitive structure. As $k$ increases, the checked attribute becomes less upwardly biased. Therefore, the probability of purchase decreases, but the equilibrium can be sustained at higher prices.
The top equilibria are extremal within the set of all equilibria, in a sense that we make precise using Lemma 1. Therefore, by analyzing the top equilibria, we can draw conclusions about all equilibria. Theorem 2 shows that the sender can strictly benefit from communication if and only if the sender strictly prefers one of the top equilibria to the no-communication outcome. The key condition is that the top-$(N-1)$ equilibrium exists. Theorem 3 shows that the sender-preferred top-$k$ equilibrium is Pareto efficient within the set of all equilibria. Moreover, for each $k$, there is some price at which the top-$k$ equilibrium is sender-optimal among all equilibria.
Our main model assumes that the price is exogenous. Next, we allow the sender to choose the price (before observing the state). In this case, Theorem 4 shows that the sender-optimal outcome is either one of the top equilibria or else the no-communication outcome.
Finally, we consider an extension in which the receiver can check $n$ of the $N$ attributes, where $1<n<N$. We construct a natural analogue of each top-$k$ equilibrium. The sender’s message reveals which $k$ attributes are best, and the receiver checks attributes in that set. In Theorem 5, we characterize the interval of prices at which the top-$k$ equilibrium exists. As $k$ increases, this interval shifts upward.
The rest of the paper is organized as follows. After discussing related literature, we present the model in Section 2. In Section 3, we characterize whether the top equilibria exist. In Section 4, we analyze the optimality of the top equilibria, and we characterize whether the sender benefits from communication. In Section 5, we consider the extension in which the receiver can check multiple attributes. Section 6 is the conclusion. Proofs are in Appendix A.
@@
The verification structure in our model follows Glazer and Rubinstein (2004). In that paper, the receiver can check one component of the multidimensional state before choosing a binary action. [^1] Glazer and Rubinstein (2006) study a complementary problem where the sender chooses which hard evidence to present to the receiver. In a more general evidence setting, Hart et al. (2017) establish an equivalence between equilibria (without commitment) and receiver-optimal mechanisms (with commitment). The sender has a strict, state-independent preference for one action over the other. Crucially, the receiver has commitment power. Their paper solves for the receiver-optimal mechanism. Then they show that the receiver-optimal mechanism can be supported as an equilibrium, without receiver commitment. By contrast, we study the sender’s gains from communication, without receiver commitment. Carroll and Egorov (2019) extend the setting of Glazer and Rubinstein (2004) by allowing the sender to have a general utility function over the receiver’s induced beliefs. They give a necessary and sufficient condition on the utility function for the receiver to perfectly learn the state. Their condition is not satisfied in our setting.
@@
In many other models of verification, the receiver observes an exogenous signal about the state. [^2] An exception is Hancart (2024). The receiver chooses from an arbitrary set of Blackwell experiments about the state. That paper characterizes whether it is optimal for the receiver to choose a Blackwell-dominated experiment. Kattwinkel (2019) and Silva (2024) give conditions under which the receiver, with commitment power, can benefit from communication, i.e., screening the sender. Weksler and Zik (2025) give conditions under which communication can influence the receiver’s action, when the sender has private information about the informativeness of the receiver’s signal. In these papers, the sender’s message influences the sensitivity of the receiver’s action to his signal realization. In our model, by contrast, the sender’s message influences which signal the receiver observes. Finally, communication equilibria have been analyzed in models where the receiver can (a) detect lies with an exogenous probability (Balbuzanov, 2019; Dziuda and Salas, 2018), (b) choose to detect lies at a cost (Sadakane and Tam, 2023), or (c) choose to learn the state at a cost (Bijkerk et al., 2018; Venkatesh, 2024).
@@
In our model, the receiver can verify the state, but the sender’s message is costless and unrestricted. An alternative modeling approach is taken in disclosure games (Milgrom, 1981; Grossman, 1981). In that model, the sender can “prove” to the receiver any true statement about the state. Thus, messages have intrinsic meaning, and the true state determines which messages are feasible. Subsequent work imposes structure on the class of true statements that can be proven. In Fishman and Hagerty (1990), the sender can disclose one dimension of a multidimensional state. In Dziuda (2011), the sender can present verifiable “arguments,” and the number of arguments is unknown to the receiver.
@@ 2 Model
The receiver has a limited capacity to learn about the product. Following Glazer and Rubinstein (2004), we assume that the receiver can check exactly one component of the state $\theta$ [^4] In Section 5, we consider the case in which the receiver can check multiple components. Checking component $i$ perfectly reveals $\theta_{i}$. The timing is as follows. The sender observes the state and sends a costless message to the receiver. The receiver sees the message and then chooses which of the $N$ dimensions to check. After observing whether the checked component is good (11 11) or bad (00 0), the receiver chooses whether to buy the product.
@@ 3.2 Top equilibria
We now construct a family of influential equilibria, indexed by $k=1,\ldots,N-1$. In the top-$k$ equilibrium, the sender’s message reveals which $k$ attributes are best, with ties broken uniformly. The receiver randomly selects one of these $k$ “recommended” attributes to check. Then the receiver buys if and only if the checked attribute is good. For example, suppose that $N=3$. Consider the top-22 22 equilibrium. If $\theta=(1,0,0)$, then with probability $1/2$ the sender’s message indicates that “attributes 1 and 2 are the two best attributes,” and with probability $1/2$ the sender’s message indicates that “attributes 1 and 3 are the two best attributes.” To be sure, messages have no intrinsic meaning; their interpretation is determined endogenously by the sender’s messaging strategy.
Formally, for $k=1,\ldots,N-1$, the top-$k$ strategy profile $(m_{k};c_{k},b_{k})$ is defined as follows. Let $\mathcal{P}_{k}$ denote the collection of all $k$-element subsets of $[N]$. For each $A\in\mathcal{P}_{k}$, let $\lvert\theta_{A}\rvert=\textstyle\sum_{i\in A}\theta_{i}$. That is, $\lvert\theta_{A}\rvert$ counts the number of attributes in $A$ that are good. Set $M=\mathcal{P}_{k}$. [^6] Technically, the set $M$ is fixed. We identify $\mathcal{P}_{k}$ with some subset of $M$. We assume that the receiver treats every message $m\in M\setminus\mathcal{P}_{k}$ as if he had received the same fixed message $m_{0}\in\mathcal{P}_{k}$. For each state $\theta$, let $m_{k}(\theta)$ be the uniform distribution over $\operatorname*{argmax}\nolimits_{A\in\mathcal{P}_{k}}\,\lvert\theta_{A}\rvert$. For each $A\in\mathcal{P}_{k}$, let $c_{k}(A)$ be the uniform distribution over $A$. Finally, let $b_{k}(A,i,z)=z$ for each $(A,i,z)\in\mathcal{P}_{k}\times[N]\times\{0,1\}$. That is, the receiver buys if and only if the attribute that he checks is good. If the top-$k$ strategy profile is an equilibrium, then we call it the top-$k$ equilibrium, and we say that the top-$k$ equilibrium exists.
Our first result gives a necessary and sufficient condition for the existence of the top-$k$ equilibrium. To state this condition, we introduce notation for the signal that the receiver obtains under the top-$k$ strategy profile. For each $k\in[N]$, let $T^{k}$ ($B^{k}$) denote a uniform draw from the top (bottom) $k$ attributes, where these top (bottom) attributes are determined by uniform tie-breaking. In state $\theta$, we know that among the top $k$ attributes, $\lvert\theta\rvert\wedge k$ are good; among the bottom $k$ attributes, $(\lvert\theta\rvert-(N-k))_{+}$ are good. Therefore, given the state $\theta$, the random variables $T^{k}$ and $B^{k}$ are conditionally independent Bernoulli random variables, with
@@
Figure 1 illustrates these price thresholds and the associated buying probabilities in a numerical example. In this example, $N=4$ and $v(\theta)=\lvert\theta\rvert$. The state distribution is such that $\lvert\theta\rvert$ is equally likely to take the values $0,\ldots,4$. For each $k=1,2,3$, we plot the buying probability $\mathbf{P}(T^{k}=1)$ over the range of prices at which the top-$k$ equilibrium exists. The price range $[\underaccent{\bar}{p}_{k},\bar{p}_{k}]$ shifts up as $k$ increases. Intuitively, as $k$ increases, the recommended attributes become less upwardly biased relative to a randomly chosen attribute. The dashed line indicates the buying probability without communication: for $p<\underaccent{\bar}{\nu}$, the receiver always buys; for $\underaccent{\bar}{\nu}<p<\bar{\nu}$, the receiver buys if and only if the checked attribute is good. [^7] For $p>\bar{\nu}$, the receiver never buys. This part of the graph is omitted from the plot. If $\underaccent{\bar}{\nu}<p\leq\bar{p}_{3}$, then there exists a top equilibrium that the sender strictly prefers to the no-communication outcome.
@@ 4.1 Equilibrium bounds
Lemma 1 uses only the sender’s incentive constraints, so the bound in (1) is independent of the price $p$ and the receiver’s utility function $v$. The sender’s incentive constraints limit how steeply the buying probability can increase as a function of the number $\lvert\theta\rvert$ of good attributes. The precision of the verification technology determines the maximal steepness. Indeed, without verification, the buying probability must be flat. If the receiver could perfectly verify the entire state, then the buying probability could be arbitrarily steep.
To prove Lemma 1, we fix an arbitrary equilibrium and then analyze a special class of deviations. Consider some type $\theta$ with $\lvert\theta\rvert=j>1$. Choose a dimension $i$ such that $\theta_{i}=1$, and consider the state $\theta^{\prime}=\theta-e_{i}$, where $e_{i}$ is the $i$-th standard basis vector. Type $\theta^{\prime}$ can mimic type $\theta$ by sending some message that is sent by type $\theta$ in equilibrium. If the receiver checks any dimension other than dimension $i$, then he cannot distinguish type $\theta$ from type $\theta^{\prime}$, so the outcome must be the same as if type $\theta$ had sent that message. There are $j$ feasible choices for this dimension $i$. Averaging over these $j$ choices, we conclude that the average buying probability over those $j$ states is at least $(j-1)/j$ times the buying probability in state $\theta$. By further averaging over all states $\theta$ with $j$ good attributes, we obtain the inequality in (1).
@@
The vector $q^{k}$ satisfies (1) with equality for each $j$ such that $q_{j-1}^{k}<1$. Let $Q$ be the set of all vectors $q$ in $[0,1]^{N+1}$ that satisfy (1). The set $Q$ is a polytope. It is easily verified that each vector $q^{k}$ is an extreme point of $Q$. Let $Q_{0}$ be the set of all buying probability vectors that are induced by some equilibrium. The set $Q_{0}$, unlike $Q$, depends on the price $p$ and the utility function $v$. This set $Q_{0}$ is difficult to characterize, but we know from Lemma 1 that $Q_{0}$ is a subset of $Q$. It follows that the vector $q^{k}$ is an extreme point of $Q_{0}$ whenever the top-$k$ equilibrium exists.

### v3

@@ 1 Introduction
Theorem˜1 characterizes whether each top-$k$ equilibrium exists. This family of equilibria has an intuitive structure. As $k$ increases, the checked attribute becomes less upwardly biased. Therefore, the probability of purchase decreases, but the equilibrium can be sustained at higher prices.
Lemma˜1 derives tight bounds on equilibrium buying probabilities using only the sender’s incentive constraints: the buying probability cannot increase more than proportionally with the number of good attributes. Crucially, the top-$k$ buying probability vectors satisfy these bounds with equality—making the top equilibria extreme points of the feasible set of all equilibrium buying probability vectors. Therefore, by analyzing the top equilibria, we can draw conclusions about all equilibria.
Theorem˜2 shows that the sender can strictly benefit from communication if and only if the sender strictly prefers one of the top equilibria to the no-communication outcome. The non-trivial direction says that when no top equilibrium benefits the sender, no equilibrium can. Crucially, Lemma˜1 gives bounds on the set of feasible equilibrium outcomes. If all top equilibria fail to benefit the sender, then we show that the sender cannot benefit from any outcome that satisfies these bounds and the receiver’s participation constraint.
Theorem˜3 shows that the sender-preferred top-$k$ equilibrium is Pareto efficient within the set of all equilibria. Using the extremality from Lemma˜1, we show that no feasible equilibrium buying probability vector can simultaneously give both players a higher payoff. At the highest price at which the top-$k$ equilibrium exists, the top-$k$ equilibrium is also sender-optimal among all equilibria. At this price, the top-$k$ equilibrium gives the receiver his lowest possible payoff—equal to what he would obtain by checking a randomly chosen attribute. Moreover, for each $k$, there is some price at which the top-$k$ equilibrium is sender-optimal among all equilibria.
Our main model assumes that the price is exogenous. Next, we allow the sender to choose the price (before observing the state). In this case, Theorem˜4 shows that the sender-optimal outcome is either one of the top equilibria or else the no-communication outcome. The key step is showing that the sender’s maximum attainable revenue is convex in price between any two consecutive top-equilibrium price thresholds. Therefore, its global maximum must be attained at some price threshold. At such a threshold, the associated top-$k$ equilibrium yields the maximum revenue.
Finally, we consider an extension in which the receiver can check $n$ of the $N$ attributes, where $1<n<N$. We construct a natural analogue of each top-$k$ equilibrium. The sender’s message reveals which $k$ attributes are best, and the receiver checks attributes in that set. In Theorem˜5, we characterize the interval of prices at which the top-$k$ equilibrium exists. As $k$ increases, this interval shifts upward.
The rest of the paper is organized as follows. After discussing related literature, we present the model in Section˜2. In Section˜3, we characterize whether the top equilibria exist. In Section˜4, we analyze the optimality of the top equilibria, and we characterize whether the sender benefits from communication. In Section˜5, we consider the extension in which the receiver can check multiple attributes. Section 6 is the conclusion. Proofs are in Appendix˜A.
@@
None of these results is useful in a binary setting. When the receiver’s action is binary and the sender has nontrivial state-independent preferences, cheap talk cannot be influential: all equilibrium messages must induce the same (mixed) action. In our model, unlike in Chakraborty and Harbaugh (2007, 2010), revealing information about the ranking of the attributes cannot, by itself, change the purchase decision. Rather, this ordinal information changes the receiver’s interpretation of the information that he acquires. Thus, the verification technology in our model is precisely what enables cheap talk to be influential in our binary setting.
The broader literature on state verification was initiated by Townsend (1979), who studies optimal contracts when the principal can verify the agent’s privately observed state (realized endowment or project return) at a cost; see also Gale and Hellwig (1985) and Border and Sobel (1987). More recent work studies multi-agent allocation problems with verification but no monetary transfers. Each agent privately knows the principal’s value from allocating to that agent. Ben-Porath et al. (2014) characterize optimal mechanisms when there is a single item to allocate and verifying each agent’s information is costly. In Erlanson and Kleiner (2026), there are multiple items, and there is a cap on the number of agents that the principal can verify. These papers characterize the principal’s optimal mechanism. In our model, the receiver does not have commitment power, and we analyze the set of cheap-talk equilibria. This requires different techniques because we cannot apply a revelation principle.
@@
The verification structure in our model follows Glazer and Rubinstein (2004). In that paper, the receiver can check one component of the multidimensional state before choosing a binary action. The sender has a strict, state-independent preference for one action over the other. The receiver has commitment power, and they solve for the receiver-optimal mechanism. Then they show that the receiver-optimal mechanism can be supported as an equilibrium, without receiver commitment. [^1] Glazer and Rubinstein (2006) solve for the receiver-optimal mechanism in a complementary problem in which the sender chooses which hard evidence to present to the receiver. They show that neither commitment nor randomization has value to the receiver. Sher (2011) shows that this result extends beyond the binary setting, provided that the receiver’s payoff is a concave transformation of the sender’s payoff. More generally, Sher (2014) presents a linear program that characterizes the “dynamic” persuasion problem from Glazer and Rubinstein (2004) and then, with an additional integer constraint, characterizes the “static” persuasion problem from Glazer and Rubinstein (2006). By contrast, we study the sender’s gains from communication, without receiver commitment. The equilibrium analysis, unlike the mechanism design problem, cannot be reduced to solving a linear program. Carroll and Egorov (2019) extend the setting of Glazer and Rubinstein (2004) by allowing the sender to have a general utility function over the receiver’s induced beliefs. They give a necessary and sufficient condition on the utility function for the receiver to perfectly learn the state. Their condition is not satisfied in our setting.
@@
In many other models of verification, the receiver observes an exogenous signal about the state. [^2] An exception is Hancart (2025). The receiver chooses from an arbitrary set of Blackwell experiments about the state. That paper characterizes whether it is optimal for the receiver to choose a Blackwell-dominated experiment. Kattwinkel (2019) and Silva (2024) give conditions under which the receiver, with commitment power, can benefit from communication, i.e., screening the sender. Weksler and Zik (2025) give conditions under which communication can influence the receiver’s action, when the sender has private information about the informativeness of the receiver’s signal. In these papers, the sender’s message influences the sensitivity of the receiver’s action to his signal realization. In our model, by contrast, the sender’s message influences which signal the receiver observes. Finally, communication equilibria have been analyzed in models where the receiver can (a) detect lies with an exogenous probability (Balbuzanov, 2019; Dziuda and Salas, 2018), (b) choose to detect lies at a cost (Sadakane and Tam, 2026), or (c) choose to learn the state at a cost (Bijkerk et al., 2018; Venkatesh, 2024).
@@
In our model, the receiver can partially verify the state, but the sender’s message is costless and unrestricted. An alternative modelling approach is taken in disclosure games (Milgrom, 1981; Grossman, 1981). In that model, the sender can “prove” to the receiver any true statement about the state. Thus, messages have intrinsic meaning, and the true state determines which messages are feasible. [^3] In our model, the receiver randomly selects one attribute to verify upon receiving the sender’s message, which essentially induces the sender to transmit more information than is ultimately verified. Under deterministic verification, the communication situation here can be regarded as one where the sender chooses which evidence to present. Subsequent work imposes structure on the class of true statements that can be proven. In Fishman and Hagerty (1990), the sender can disclose one dimension of a multidimensional state. In Dziuda (2011), the sender can present verifiable “arguments,” and the number of arguments is unknown to the receiver. [^4] A comprehensive survey on “evidence” in game theory and mechanism design can be found in Ben-Porath et al. (2025).
@@ 2 Model
The receiver has a limited capacity to learn about the product. Following Glazer and Rubinstein (2004), we assume that the receiver can check exactly one component of the state $\theta$. [^6] In Section 5, we consider the case in which the receiver can check multiple components. Checking component $i$ perfectly reveals $\theta_{i}$. The timing is as follows. The sender observes the state and sends a costless message to the receiver. The receiver sees the message and then chooses which of the $N$ dimensions to check. After observing whether the checked component is good ($1$) or bad ($0$), the receiver chooses whether to buy the product.
@@ 3.2 Top equilibria
We now construct a family of influential equilibria, indexed by $k=1,\ldots,N-1$. In the top-$k$ equilibrium, the sender’s message reveals which $k$ attributes are best, with ties broken uniformly. The receiver randomly selects one of these $k$ “recommended” attributes to check. Then the receiver buys if and only if the checked attribute is good. For example, suppose that $N=3$. Consider the top-$2$ equilibrium. If $\theta=(1,0,0)$, then with probability $1/2$ the sender’s message indicates that “attributes 1 and 2 are the two best attributes,” and with probability $1/2$ the sender’s message indicates that “attributes 1 and 3 are the two best attributes.” To be sure, messages have no intrinsic meaning; their interpretation is determined endogenously by the players’ equilibrium strategies.
Formally, for $k=1,\ldots,N-1$, the top-$k$ strategy profile $(m_{k};c_{k},b_{k})$ is defined as follows. Let $\mathcal{P}_{k}$ denote the collection of all $k$-element subsets of $[N]$. For each $A\in\mathcal{P}_{k}$, let $\lvert\theta_{A}\rvert=\textstyle\sum_{i\in A}\theta_{i}$. That is, $\lvert\theta_{A}\rvert$ counts the number of good attributes in $A$. Set $M=\mathcal{P}_{k}$. [^8] Technically, the set $M$ is fixed. We identify $\mathcal{P}_{k}$ with some subset of $M$. We assume that the receiver treats every message $m\in M\setminus\mathcal{P}_{k}$ as if he had received the same fixed message $m_{0}\in\mathcal{P}_{k}$. For each state $\theta$, let $m_{k}(\theta)$ be the uniform distribution over $\operatorname*{argmax}\nolimits_{A\in\mathcal{P}_{k}}\,\lvert\theta_{A}\rvert$. For each $A\in\mathcal{P}_{k}$, let $c_{k}(A)$ be the uniform distribution over $A$. Finally, let $b_{k}(A,i,z)=z$ for each $(A,i,z)\in\mathcal{P}_{k}\times[N]\times\{0,1\}$. That is, the receiver buys if and only if the attribute that he checks is good. If the top-$k$ strategy profile is an equilibrium, then we call it the top-$k$ equilibrium, and we say that the top-$k$ equilibrium exists.
Our first result gives a necessary and sufficient condition for the existence of the top-$k$ equilibrium. To state this condition, we introduce notation for the signal that the receiver obtains under the top-$k$ strategy profile. For each $k\in[N]$, let $T^{k}$ ($B^{k}$) denote a uniform draw from the top (bottom) $k$ attributes, where these top (bottom) attributes are determined by uniform tie-breaking. In state $\theta$, we know that among the top $k$ attributes, $\lvert\theta\rvert\wedge k$ are good; among the bottom $k$ attributes, $(\lvert\theta\rvert-(N-k))_{+}$ are good. [^9] Throughout this paper, $\wedge$ denotes the minimum operator and $(\cdot)_{+}$ denotes the positive part operator, i.e., $\max\{\cdot,0\}$. Therefore, given the state $\theta$, the random variables $T^{k}$ and $B^{k}$ are conditionally independent Bernoulli random variables, with
@@
Figure˜1 illustrates these price thresholds and the associated buying probabilities in a numerical example. In this example, $N=4$ and $v(\theta)=\lvert\theta\rvert$. The state distribution is such that $\lvert\theta\rvert$ is equally likely to take the values $0,\ldots,4$. For each $k=1,2,3$, we plot the buying probability $\mathbf{P}(T^{k}=1)$ over the range of prices at which the top-$k$ equilibrium exists. The price range $[\underaccent{\bar}{p}_{k},\bar{p}_{k}]$ shifts up as $k$ increases. Intuitively, as $k$ increases, the recommended attributes become less upwardly biased relative to a randomly chosen attribute. The dashed line indicates the buying probability without communication: for $p<\underaccent{\bar}{\nu}$, the receiver always buys; for $\underaccent{\bar}{\nu}<p<\bar{\nu}$, the receiver buys if and only if the checked attribute is good. [^10] For $p>\bar{\nu}$, the receiver never buys. This part of the graph is omitted from the plot. If $\underaccent{\bar}{\nu}<p\leq\bar{p}_{3}$, then there exists a top equilibrium that the sender strictly prefers to the no-communication outcome.
@@ 4.1 Equilibrium bounds
Lemma˜1 uses only the sender’s incentive constraints, so the bound in (1) is independent of the price $p$ and the receiver’s utility function $v$. The sender’s incentive constraints limit how steeply the buying probability can increase as a function of the number $\lvert\theta\rvert$ of good attributes. The precision of the verification technology determines the maximal steepness. Indeed, without verification, the buying probability must be flat. If the receiver could perfectly verify the entire state, then the buying probability could be arbitrarily steep.
To prove Lemma˜1, we fix an arbitrary equilibrium and then analyze a special class of deviations. Consider some type $\theta$ with $\lvert\theta\rvert=j>1$. Choose a dimension $i$ such that $\theta_{i}=1$, and consider the state $\theta^{\prime}=\theta-e_{i}$, where $e_{i}$ is the $i$-th standard basis vector. Type $\theta^{\prime}$ can mimic type $\theta$ by sending some message that is sent by type $\theta$ in equilibrium. If the receiver checks any dimension other than dimension $i$, then he cannot distinguish type $\theta$ from type $\theta^{\prime}$, so the outcome must be the same as if type $\theta$ had sent that message. There are $j$ feasible choices for this dimension $i$. Averaging over these $j$ choices, we conclude that the average buying probability over those $j$ states is at least $(j-1)/j$ times the buying probability in state $\theta$. By further averaging over all states $\theta$ with $j$ good attributes, we obtain the inequality in (1).
@@
The vector $q^{k}$ satisfies (1) with equality for each $j$ such that $q_{j-1}^{k}<1$. Let $Q$ be the set of all vectors $q$ in $[0,1]^{N+1}$ that satisfy (1). The set $Q$ is a polytope. It is easily verified that each vector $q^{k}$ is an extreme point of $Q$. Let $Q_{0}$ be the set of all buying probability vectors that are induced by some equilibrium. The set $Q_{0}$, unlike $Q$, depends on the price $p$ and the utility function $v$. This set $Q_{0}$ is difficult to characterize, but we know from Lemma˜1 that $Q_{0}$ is a subset of $Q$. It follows that the vector $q^{k}$ is an extreme point of $Q_{0}$ whenever the top-$k$ equilibrium exists.